OpenStudy (anonymous):

solve by elimination x-7y = -34 7x-7y=14

OpenStudy (anonymous):

x = 8 y = 6

OpenStudy (anonymous):

bboy check ur answer

OpenStudy (anonymous):

(-1)x-7y = -34 7x-7y=14 -x + 7y = 34 7x - 7y = 14 6x = 48 /6 /6 x= 8 7x-7y=14 7(8) - 7y = 14 56 - 7y = 14 -56 -56 -7y = -42 /-7 /-7 y = 6 (8,6)